OpenStudy (anonymous):

Solve for n. n/8=5 and 1/5 A.13/20 B.1 and 7/13 C.41 and 3/5 D.208

OpenStudy (alexandervonhumboldt2):

n/8=5 1/5 first cnvert to improper fraction: a b/c=(a*c+b)/c so it would be n/8=26/5 n=26/5*8=41.6=41 3/5

OpenStudy (anonymous):

so it is C

OpenStudy (anonymous):

@AlexandervonHumboldt2